a survey of of 150 people indicated that 80people would buy new shoes in the folllowing year, 95 people would buy a snappy hat, and that 32people would buy neither.
a) how many people would buy both new shoes and a snappy hat this year?
b) how many people would buy new shoes or a snappy hat this year?

Answer by robertb(5830)

Let n = # of people who would buy BOTH new shoes AND a snappy hat this year. Then
150 = 32 + (80 - n) + n + (95 - n)
==> 150 = 207 - n
==> -57 = -n
==> n = 57.
Number of people who would buy new shoes OR a snappy hat this year = 150 - 32 = 118.